# Q3R: Quadratic Reweighted Rank Regularizer for Effective Low-Rank Training

### v1

@@ Appendix B Computation of Q3R value
$\bigl\langle\mathbf{W},\mathcal{R}_{\mathbf{W}^{\prime},\epsilon}(\mathbf{W})\bigr\rangle$ $=\mathrm{tr}\bigl(\mathbf{W}^{\top}\mathbf{W}\bigr)$
$\quad+\;\mathrm{tr}\!\bigl(\mathbf{U}^{\top}\mathbf{W}\,\mathbf{W}^{\top}\mathbf{U}\,\bigl(\epsilon\,\mathbf{\Sigma}^{-1}-\mathbf{I}\bigr)\bigr)$
$\quad+\;\mathrm{tr}\!\bigl(\mathbf{V}^{\top}\mathbf{W}^{\top}\mathbf{W}\,\mathbf{V}\,\bigl(\epsilon\,\mathbf{\Sigma}^{-1}-\mathbf{I}\bigr)\bigr)$
$\quad+\;\mathrm{tr}\!\bigl(\mathbf{V}^{\top}\mathbf{W}^{\top}\mathbf{U}\,\bigl(\epsilon\,\mathbf{\Sigma}^{-1}-\mathbf{I}\bigr)\,\mathbf{U}^{\top}\mathbf{W}\,\mathbf{V}\,\bigl(\epsilon\,\mathbf{\Sigma}^{-1}-\mathbf{I}\bigr)\bigr).$
For each iteration we calculate the quadratic regularizer $\langle\mathbf{W},\mathcal{R}_{\mathbf{W}^{\prime},\epsilon}(\mathbf{W})\rangle$ for weight matrices $\mathbf{W}$. For algorithmic simplicity, We now re-arrange the summand of our Q3R regularizer and show the simplified expression for this inner product in Algorithm˜3. We later show in Equation˜13, where we derive the expression of this inner product and we show that this matches the one proposed in Algorithm˜3.
$f$ $=\underbrace{\operatorname{\operatorname{tr}}\bigl(\mathbf{W}^{\top}\mathbf{W}\bigr)}_{t_{1}}+\underbrace{\operatorname{\operatorname{tr}}\!\bigl(\mathbf{V}^{\top}\mathbf{W}^{\top}\mathbf{U}\,\mathbf{S}\,\mathbf{U}^{\top}\mathbf{W}\,\mathbf{V}\,\mathbf{S}\bigr)}_{t_{2}}+\underbrace{\operatorname{\operatorname{tr}}\!\bigl(\mathbf{V}^{\top}\mathbf{W}^{\top}\mathbf{W}\,\mathbf{V}\,\mathbf{S}\bigr)}_{t_{3}}+\underbrace{\operatorname{\operatorname{tr}}\!\bigl(\mathbf{U}^{\top}\mathbf{W}\,\mathbf{W}^{\top}\mathbf{U}\,\mathbf{S}\bigr)}_{t_{4}},$
$\mathbf{S}$ $=\epsilon\,\mathbf{\Sigma}^{-1}-\mathbf{I},$
$\mathbf{T}$ $=\mathbf{W}\,\mathbf{V},\quad\mathbf{B}=\mathbf{W}^{\top}\mathbf{U},\quad\mathbf{C}=\mathbf{U}^{\top}\mathbf{W}\,\mathbf{V},\quad\mathbf{M}=\mathbf{S}^{1/2}\,\mathbf{C}\,\mathbf{S}^{1/2}.$
Using $\operatorname{\operatorname{tr}}(\mathbf{A}^{\top}\mathbf{A})=\|\mathbf{A}\|_{F}^{2}$ and cyclicity:
$t_{1}$ $=\operatorname{\operatorname{tr}}(\mathbf{W}^{\top}\mathbf{W})=\|\mathbf{W}\|_{F}^{2},$
$t_{2}$ $=\operatorname{\operatorname{tr}}\!\bigl(\mathbf{S}\,(\mathbf{U}^{\top}\mathbf{W}\mathbf{V})\,\mathbf{S}\,(\mathbf{U}^{\top}\mathbf{W}\mathbf{V})\bigr)=\|\mathbf{M}\|_{F}^{2},$
$t_{3}$ $=\operatorname{\operatorname{tr}}\!\bigl(\mathbf{S}\,(\mathbf{W}\mathbf{V})^{\top}(\mathbf{W}\mathbf{V})\bigr)=\|\mathbf{T}\,\mathbf{S}^{1/2}\|_{F}^{2},$
$t_{4}$ $=\operatorname{\operatorname{tr}}\!\bigl(\mathbf{S}\,(\mathbf{W}^{\top}\mathbf{U})^{\top}(\mathbf{W}^{\top}\mathbf{U})\bigr)=\|\mathbf{B}\,\mathbf{S}^{1/2}\|_{F}^{2}.$
Putting it all together, with $\mathbf{R}=\mathbf{S}^{1/2}$:
$f(\mathbf{W},\mathbf{U},\mathbf{V})=\|\mathbf{W}\|_{F}^{2}+\|\mathbf{M}\|_{F}^{2}+\|\mathbf{T}\,\mathbf{R}\|_{F}^{2}+\|\mathbf{B}\,\mathbf{R}\|_{F}^{2},\qquad\mathbf{R}=(\epsilon\,\mathbf{\Sigma}^{-1}-\mathbf{I})^{1/2}.$ (13)
@@
Algorithm 3 Computation of the Q3R  function value $\operatorname{Q3R}_{\mathbf{W}^{\prime},\epsilon}(\mathbf{W})$
1:Input: $W\in\mathbb{R}^{d_{1}\times d_{2}}$, $U\in\mathbb{R}^{d_{1}\times r}$, $V\in\mathbb{R}^{d_{2}\times r}$, $S\in\mathbb{R}^{r\times r}$
2:Output:e $f(W,U,V,S)=\|W\|_{F}^{2}+\|\,M\|_{F}^{2}+\|\,T\,R\|_{F}^{2}+\|\,B\,R\|_{F}^{2}$
3:1. Compute projections
@@
6:2. Form intermediate products
[…]
8: Compute symmetric square‐root $R$ of $S$: $R\,R^{\top}=S$ $O(r^{3})$
[…]
10:3. Evaluate Frobenius norms
@@
15:return $f\leftarrow t_{1}+t_{2}+t_{3}+t_{4}$
@@
$\boxed{T_{\mathrm{total}}=2\,d_{1}d_{2}r\;+\;(d_{1}+d_{2})r^{2}\;+\;2r^{3}\;+\;\bigl(d_{1}d_{2}+(d_{1}+d_{2})r+r^{2}\bigr)=O\bigl(d_{1}d_{2}r\;+\;(d_{1}+d_{2})r^{2}\;+\;r^{3}\bigr)}$
@@ Appendix C Computation of the Gradient of Q3R
Using $\epsilon\,\mathbf{\Sigma}^{-1}=\mathbf{I}+\mathbf{S}$, we rewrite each term:
@@
Collecting like‐terms in powers of $\mathbf{S}$ gives the compact form
$\boxed{\mathcal{R}_{\mathbf{W}^{\prime},\epsilon}(\mathbf{W})=\underbrace{\mathbf{W}}_{T_{1}}+\underbrace{\mathbf{U}\,\mathbf{S}\,\mathbf{U}^{\top}\,\mathbf{W}}_{T_{2}}+\underbrace{\mathbf{W}\,\mathbf{V}\,\mathbf{S}\,\mathbf{V}^{\top}}_{T_{3}}+\underbrace{\mathbf{U}\,\mathbf{S}\,\mathbf{U}^{\top}\,\mathbf{W}\,\mathbf{V}\,\mathbf{S}\,\mathbf{V}^{\top}}_{T_{4}}}$
@@
Algorithm 4 Computation gradient of Q3R : Compute $\mathcal{R}_{\mathbf{W}^{\prime},\epsilon}(\mathbf{W})$
1:Input: $\mathbf{W}\!\in\!\operatorname{\mathbb{R}}^{d_{1}\times d_{2}}$, $\mathbf{U}\!\in\!\operatorname{\mathbb{R}}^{d_{1}\times r}$, $\mathbf{V}\!\in\!\operatorname{\mathbb{R}}^{d_{2}\times r}$, singular values $\boldsymbol{\sigma}\!\in\!\operatorname{\mathbb{R}}^{r}_{>0}$, threshold $\epsilon$
2:Output: $\mathbf{G}=\mathcal{R}_{\mathbf{W}^{\prime},\epsilon}(\mathbf{W})$
3:
@@
[…]
7:
[…]
10:
[…]
13:for $i=1,\dots,r$ do
14:   for $j=1,\dots,r$ do
15:   $E_{ij}\leftarrow(S_{\text{shift}})_{ii}\cdot C_{ij}\cdot(S_{\text{shift}})_{jj}$
[…]
18:
@@
25:return $\mathbf{gradient}$
@@ D.1 Experimental Protocol
In the experiments, we compared the training of unregularized models, as well as models regularized by Q3R , LoRITa [3] and LoRA [28]. Due to the limitations of some techniques not providing strong truncation guide lines, we truncate each method at various truncation ranks $r$. Ensuring such that $r$ is chosen to always ensure that the factor matrix pairs always have less original parameters, to minimize a practical environment. In Table˜1 we find that running with $p=.20$ or $r=19$ and $\lambda=0.001$ performs best given the truncation and accuracy trade-offs.
@@
We fine-tuned the pre-trained RoBERTa-Base model (from Hugging Face) on all nine GLUE tasks using a maximum sequence length of 512 and a batch size of 16 (30 for CoLA). We run all fine-tuning experiments with a target rank as low as $4$. The best performing setup had ’reweighting period’ of $3$ and $\lambda\in\{1.5,2\}$. As shown in Table˜3, we compare Q3R with LoRA and dense fine-tuning. The tasks and their metrics are summarized below:
@@ D.2 Fine-tuning Experiments
To assess the potential of Q3R for fine-tuning large language models (LLMs), we conducted an additional experiment on Llama 3.2–3B using an NVIDIA A5000 GPU. Since the original meta-llama/Llama-3.2-3B checkpoint provides only the pre-trained language model parameters, we instantiated the LlamaForSequenceClassification module, which attaches an additional linear projection, commonly referred to as the classification head, on top of the final hidden-state representation. This head, whose weight matrix is registered as score.weight (and optionally score.bias), is absent from the checkpoint and was therefore initialized with random Gaussian values.
@@
To further compare our proposed method, we conducted experiemnts on Llama3.2-1B on a subset of the GLUE tasks. Table˜4 provides a comparison of our method’s performance with dense fine-tuning and LoRa.
@@
D.3 Additional pretraining experiments
Table˜5 and Table˜6 demonstrate the low rank induction techniques on ViT models. We train ViT-B on CIFAR-100 in Table˜6 and ViT-T on CIFAR-10 in Table˜5 respectively. A visual representation of the results in Table˜5 is also illustrated in Figure˜3 and that of Table˜6 in Figure˜2.
@@ D.4 Computational Aspects
For few experiments like Q3R, we used NVIDIA A50005000 to train the ViT models. The rest of the experiments were performed on NVIDIA V100100 with 3232GB memory. The fine-tuning experiemtns were all performed in NVIDIA A50005000 GPUs.
@@ D.5 Computational Overhead of Methodology
Based on the theoretical computational overhead outlined in Algorithms˜3 and 1, along with the results above, we expect Q3R to incur additional computational and memory overhead in our suboptimal implementation of Q3R. We attribute the lower Reserved GPU Memory from implementation differences in the optimizer.
@@ E.3 Merits of Low-Rank Initialization
In our experiments, we implemented spectral initialization where the regularized weight matrix has a rank greater than or equal to the target rank hyperparameter specified in Q3R. Our leading hypothesis for the poor validation accuracy following initialization is that Q3R imposes a strong constraint on the optimization landscape when the rank is low, thereby limiting the model’s capacity to explore more expressive solutions during training. This hypothesis is further supported by the finetuning results on RoBERTa in Table˜3 where Q3R is applied to the pretrained model and continuously tuned on the GLUE benchmark tasks. We achieve notable results that support the general intuition that models to tend to learn better representation through expressive architectures and training regimes. We experimented on only the smaller data to study how the initialization affects the performance of Q3R. We provide one example of the empirical evidence performed on CIFAR-10 on ViT-Tiny with $r=0.05$ and $\lambda=0.001$ in Table˜12. We observe that Lowrank initialization is unable to surpass the performance of Q3R without such constraint on the initialization as reported in Table˜12. However, the accuracy lies within $1\%$ range of the fullrank initialized model which proves that our proposed method can be implemented in resource constrained setup as well.
Table 12: Comparison of AdamQ3R with and without LowRank initialization.
Parameter retention AdamQ3R + LowRank AdamQ3R (Q3R, no LowRank)

### v2

@@ Appendix B Computation of Regularizer Value of Q3R
$\operatorname{Q3R}_{\mathbf{W}^{\prime},\epsilon}(\mathbf{W})=\frac{1}{2}\bigl\langle\mathbf{W},\mathcal{R}_{\mathbf{W}^{\prime},\epsilon}(\mathbf{W})\bigr\rangle$ $=\frac{1}{2}\Big(\operatorname{tr}\bigl(\mathbf{W}^{\top}\mathbf{W}\bigr)$
$\quad+\;\operatorname{tr}\bigl(\mathbf{U}^{\top}\mathbf{W}\,\mathbf{W}^{\top}\mathbf{U}\,\bigl(\epsilon\,\mathbf{\Sigma}^{-1}-\mathbf{I}\bigr)\bigr)$
$\quad+\;\operatorname{tr}\bigl(\mathbf{V}^{\top}\mathbf{W}^{\top}\mathbf{W}\,\mathbf{V}\,\bigl(\epsilon\,\mathbf{\Sigma}^{-1}-\mathbf{I}\bigr)\bigr)$
$\quad+\;\operatorname{tr}\bigl(\mathbf{V}^{\top}\mathbf{W}^{\top}\mathbf{U}\,\bigl(\epsilon\,\mathbf{\Sigma}^{-1}-\mathbf{I}\bigr)\,\mathbf{U}^{\top}\mathbf{W}\,\mathbf{V}\,\bigl(\epsilon\,\mathbf{\Sigma}^{-1}-\mathbf{I}\bigr)\bigr)\Big).$
For algorithmic simplicity, we now re-arrange the terms, leading to a simplified expression for this inner product in Algorithm˜3: Rewriting these terms, we obtain
$\operatorname{Q3R}_{\mathbf{W}^{\prime},\epsilon}(\mathbf{W})=$
$=\frac{1}{2}\Big(\underbrace{\operatorname{\operatorname{tr}}\bigl(\mathbf{W}^{\top}\mathbf{W}\bigr)}_{t_{1}}+\underbrace{\operatorname{\operatorname{tr}}\!\bigl(\mathbf{V}^{\top}\mathbf{W}^{\top}\mathbf{U}\,\mathbf{S}\,\mathbf{U}^{\top}\mathbf{W}\,\mathbf{V}\,\mathbf{S}\bigr)}_{t_{2}}+\underbrace{\operatorname{\operatorname{tr}}\!\bigl(\mathbf{V}^{\top}\mathbf{W}^{\top}\mathbf{W}\,\mathbf{V}\,\mathbf{S}\bigr)}_{t_{3}}+\underbrace{\operatorname{\operatorname{tr}}\!\bigl(\mathbf{U}^{\top}\mathbf{W}\,\mathbf{W}^{\top}\mathbf{U}\,\mathbf{S}\bigr)}_{t_{4}}\Big),$
where $\mathbf{S}=\epsilon\,\mathbf{\Sigma}^{-1}-\mathbf{I}=\operatorname{diag}(\mathbf{s}_{\text{shift}})$, $\mathbf{A}=\mathbf{U}^{\top}\mathbf{W}$, $\mathbf{B}=\mathbf{W}\mathbf{V}$ and $\mathbf{C}=\mathbf{U}^{\top}\mathbf{W}\,\mathbf{V}$. Using $\operatorname{\operatorname{tr}}(\mathbf{A}^{\top}\mathbf{A})=\|\mathbf{A}\|_{F}^{2}$ and cyclicity, we see that the $t_{i}$ can be computed for $t=1,2,3,4$ such that
$t_{1}$ $=\|\mathbf{W}\|_{F}^{2}=\sum_{i=1}^{d_{1}}\sum_{j=1}^{d_{2}}\mathbf{W}_{ij}^{2},$
$t_{2}$ $=\operatorname{tr}(\mathbf{C}^{\top}\mathbf{S}\mathbf{C}\mathbf{S})=\sum_{i=1}^{r}\sum_{j=1}^{r}(\mathbf{s}_{\text{shift}})_{i}(\mathbf{s}_{\text{shift}})_{j}\mathbf{C}_{ij}^{2},$
$t_{3}$ $=\operatorname{tr}\left(\mathbf{B}^{\top}\mathbf{B}\mathbf{S}\right)=\sum_{i=1}^{d_{1}}\sum_{j=1}^{r}\mathbf{B}_{ij}^{2}(\mathbf{s}_{\text{shift}})_{j},$
$t_{4}$ $=\operatorname{tr}\bigl(\mathbf{A}\mathbf{A}^{\top}\mathbf{S}\bigr)=\sum_{i=1}^{r}\sum_{j=1}^{d_{2}}(\mathbf{s}_{\text{shift}})_{i}\mathbf{A}_{ij}^{2}.$
The outline of the resulting algorithm is as follows.
@@
Algorithm 3 Computation of Q3R function value $\operatorname{Q3R}_{\mathbf{W}^{\prime},\epsilon}(\mathbf{W})$
@@
$1\mathbf{s}_{\text{shift}}\leftarrow\epsilon/\max(\boldsymbol{\sigma},\epsilon)$-1  $\triangleright\;2r$ multiplications and subtractions
@@
[…]
9:return $f\leftarrow(t_{1}+t_{2}+t_{3}+t_{4})/2$  $\triangleright\;3$ additions and 11 multiplication
In Algorithm˜3, we denote by $\mathbf{W}^{\circ 2}$ the matrix of entrywise squares of $\mathbf{W}$, by $\mathbf{s}*\mathbf{M}$ the row-wise scalar multplication of $\mathbf{s}$ with $\mathbf{M}$ and by $\operatorname{sum}(\cdot,1)$ and $\operatorname{sum}(\cdot,2)$ row- and column-wise summation of matrix entries, respectively.
@@
$\text{FLOPs}=4rd_{1}d_{2}+2r^{2}d_{2}+2rd_{2}+3rd_{1}+4r^{2}+2r=O(rd_{1}d_{2})$
@@ Appendix C Computation of the Gradient of Q3R
In this section, we outline a simple algorithm that, given reference weight matrix $\mathbf{W}^{\prime}\in\operatorname{\mathbb{R}}^{d_{1}\times d_{2}}$, current layer weight matrix $\mathbf{W}\in\operatorname{\mathbb{R}}^{d_{1}\times d_{2}}$ and smoothing parameter $\epsilon>0$, computes the gradient $\nabla_{\mathbf{W}}\operatorname{Q3R}_{\mathbf{W}^{\prime},\epsilon}(\mathbf{W})$ of the quadratic reweighted rank regularizer $\operatorname{Q3R}_{\mathbf{W}^{\prime},\epsilon}(\mathbf{W})$ defined in (6) of Section 4.1.
As mentioned in Section 4.2 and after the proof of Lemma˜4.1.2, it follows from the self-adjointness of the reweighting operator $\mathcal{R}_{\mathbf{W}^{\prime},\epsilon}(\cdot)$ that
$\nabla_{\mathbf{W}}\operatorname{Q3R}_{\mathbf{W}^{\prime},\epsilon}(\mathbf{W})=\nabla_{\mathbf{W}}\left(\frac{1}{2}\langle\mathbf{W},\mathcal{R}_{\mathbf{W}^{\prime},\epsilon}(\mathbf{W})\rangle\right)=\mathcal{R}_{\mathbf{W}^{\prime},\epsilon}(\mathbf{W})$
@@
Using the equation $\epsilon\,\mathbf{\Sigma}^{-1}=\mathbf{I}+\mathbf{S}$ for the diagonal matrix $\mathbf{S}=\epsilon\,\mathbf{\Sigma}^{-1}-\mathbf{I}$, we can rewrite each term such that
@@
Collecting like terms in powers of $\mathbf{S}$ gives the compact form
$\mathcal{R}_{\mathbf{W}^{\prime},\epsilon}(\mathbf{W})=\mathbf{W}+\underbrace{\mathbf{U}\,\mathbf{S}\,(\mathbf{U}^{\top}\,\mathbf{W}+\mathbf{U}^{\top}\,\mathbf{W}\,\mathbf{V}\,\mathbf{S}\,\mathbf{V}^{\top})}_{T_{2}}+\underbrace{\mathbf{W}\,\mathbf{V}\,\mathbf{S}\,\mathbf{V}^{\top}}_{T_{3}}$
@@
Algorithm 4 Computation of Q3R Gradient : Compute $\nabla_{\mathbf{W}}\operatorname{Q3R}_{\mathbf{W}^{\prime},\epsilon}(\mathbf{W})=\mathcal{R}_{\mathbf{W}^{\prime},\epsilon}(\mathbf{W})$
@@
$1\mathbf{s}_{\text{shift}}\leftarrow\epsilon/\max(\boldsymbol{\sigma},\epsilon)$-1  $\triangleright\;2r$ multiplications and subtractions
@@
[…]
10:return $\mathbf{grad}$
The total number of floating point operations incurred by Algorithm˜4 for computation of regularizer gradients $\nabla_{\mathbf{W}}\operatorname{Q3R}_{\mathbf{W}^{\prime},\epsilon}(\mathbf{W})$ is
$\text{FLOPs}=8rd_{1}d_{2}-2d_{1}d_{2}+4r^{2}d_{2}+r(d_{1}-d_{2})+2r=O(rd_{1}d_{2}).$
@@ D.1 Experimental Protocol
In the experiments, we compared unregularized training with models regularized by Q3R, LoRITa [1], and LoRA [27]. Due to the limitations of some techniques in providing clear truncation guidelines, we truncate each method at a range of truncation ranks $r$. We choose $r$ so that the resulting factor matrix pair has fewer parameters than the original matrix, ensuring a meaningful compression regime. In Table˜1, we find that $p=0.20$ (corresponding to $r=19$) and $\lambda=0.001$ performs best in terms of truncation–accuracy trade-offs.
@@
We fine-tuned the pretrained RoBERTa-Base model (from Hugging Face) on all nine GLUE tasks using a maximum sequence length of 512 and a batch size of 16 (30 for CoLA). We run all fine-tuning experiments with a target rank as low as $4$. The best performing setup had ’reweighting period’ of $3$ and $\lambda\in\{1.5,2\}$. As shown in Table˜3, we compare Q3R with LoRA and dense fine-tuning. The tasks and their metrics are summarized below:
@@ D.2 Fine-tuning Experiments
To assess the potential of Q3R for fine-tuning large language models (LLMs), we conducted an additional experiment on Llama 3.2–3B using an NVIDIA A5000 GPU. Since the original meta-llama/Llama-3.2-3B checkpoint provides only the pretrained language model parameters, we instantiated the LlamaForSequenceClassification module, which attaches an additional linear projection, commonly referred to as the classification head, on top of the final hidden-state representation. This head, whose weight matrix is registered as score.weight (and optionally score.bias), is absent from the checkpoint and was therefore initialized with random Gaussian values.
@@
To further compare our proposed method, we conducted experiments on Llama3.2-1B on a subset of the GLUE tasks. Table˜4 provides a comparison of our method’s performance with dense fine-tuning and LoRA.
@@
D.3 Additional pre-training experiments
Table˜5 and Table˜6 demonstrate the low-rank induction techniques on ViT models. We train ViT-B on CIFAR-100 in Table˜6 and ViT-T on CIFAR-10 in Table˜5 respectively. A visual representation of the results in Table˜5 is also illustrated in Figure˜3 and that of Table˜6 in Figure˜2.
@@ D.4 Computational Aspects
For some experiments (e.g., Q3R), we used an NVIDIA A50005000 GPU to train the ViT models. The remainder of the experiments were performed on an NVIDIA V100100 GPU with 3232GB memory. The fine-tuning experiments were all performed on NVIDIA A50005000 GPUs.
@@ D.5 Computational Overhead of Methodology
Based on the theoretical computational overhead outlined in Algorithms˜3 and 1, along with the results above, we expect Q3R to incur additional computational and memory overhead in our (currently suboptimal) implementation. We attribute the lower reserved GPU memory to implementation differences in the optimizer.
@@ E.3 Merits of Low-Rank Initialization
In our experiments, we implemented spectral initialization where the regularized weight matrix has a rank greater than or equal to the target rank hyperparameter specified in Q3R. Our leading hypothesis for the poor validation accuracy following initialization is that Q3R imposes a strong constraint on the optimization landscape when the rank is low, thereby limiting the model’s capacity to explore more expressive solutions during training. This hypothesis is further supported by the fine-tuning results on RoBERTa in Table˜3, where Q3R is applied to the pretrained model and continuously tuned on the GLUE benchmark tasks. We achieve notable results that support the general intuition that models tend to learn better representations through expressive architectures and training regimes. We experimented only on the smaller dataset to study how the initialization affects the performance of Q3R. We provide one example of empirical evidence on CIFAR-10 with ViT-Tiny, with $r=0.05$ and $\lambda=0.001$, in Table˜12. We observe that low-rank initialization is unable to surpass the performance of Q3R without such a constraint on the initialization, as reported in Table˜12. However, the accuracy lies within a $1\%$ range of the full-rank initialized model, which shows that our proposed method can be implemented in a resource-constrained setup as well.
Table 12: Comparison of AdamQ3R with and without low-rank initialization.
Parameter retention AdamQ3R + low-rank init. AdamQ3R (Q3R, no low-rank init.)